You are considering buying common stock in Grow On, Inc. The firm yesterday paid a dividend of $5.20. You have projected that dividends will grow at a rate of 10.0% per year indefinitely. The firm's beta is 2.30, the risk-free rate is 7.7%, and the market return is 10.4%. What is the most you should pay for the stock now?

$146.29

$132.99

$37.38

$41.12

$159.83
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

=7.7+2.3*(10.4-7.7)

=13.91%

Current price=D1/(Required return-Growth rate)

=(5.2*1.1)/(0.1391-0.1)

=$146.29(Approx).